if a steel ball is dropped from the lowest side of the tower, and falls 183 feet, how long will it take to hit the ground? round to the nearest tenth of a second.
Answer by Alan3354(69443)

if a steel ball is dropped from the lowest side of the tower, and falls 183 feet, how long will it take to hit the ground? round to the nearest tenth of a second.
-------------
h(t) = -16t^2 + 183 = 0
16t^2 = 183
t = sqrt(183/16)
t =~ 3.4 seconds